# Shortcuts and Identifiability in Concept-based Models from a Neuro-Symbolic Lens

### v3

@@ 1 Introduction
This raises the question of when CBMs can acquire “high-quality” concepts and inference layers. While existing studies focus on concept quality [49, 99, 55], they neglect the role of the inference layer altogether. In contrast, we cast the question in terms of whether it is possible to identify from data concepts and inference layers with the intended semantics, defined formally in Section˜3.1. We proceed to answer this question by building a novel connection with reasoning shortcuts (RSs), a well-known issue in Neuro-Symbolic (NeSy) AI whereby models achieve high accuracy by learning low-quality concepts even if the inference layer is fixed [9, 56, 84, 91]. For instance, a NeSy model for autonomous driving whose inference layer encodes the traffic laws might confuse pedestrians with red lights, as both entail the same prediction (the car has to stop) [4]. We generalize RSs to CBMs in which the inference layer is learned, and concept supervision may be absent. Our analysis shows that shortcuts can be exponentially many, even more than RSs, (we count them explicitly in Section˜B.5) and that maximum likelihood training is insufficient for attaining intended semantics. This hinders both interpretability and OOD behavior. On the positive side, we also specify conditions under which (under suitable assumptions) CBMs cannot be fooled by RSs, proving that the ground-truth concepts and inference layer can be identified (see Theorem˜3.9).
@@ 3 Reasoning Shortcuts in CBMs
Here, $\bm{\mathrm{\omega}}_{\bm{\mathrm{y}}}$ is the conditional probability of the ground-truth labels $\bm{\mathrm{y}}$. Notice that under the above assumptions CBMs attaining maximum likelihood perfectly model the ground-truth data distribution $p^{*}(\bm{\mathrm{X}},\bm{\mathrm{Y}})$, see Lemma˜C.1.
@@ 3.1 Intended Semantics and Joint Reasoning Shortcuts
Intended semantics holds if the learned concepts $\bm{\mathrm{C}}$ match the ground-truth concepts $\bm{\mathrm{G}}$ modulo simple invertible transformations – like reordering and negation (Eq.˜5, left) – and the inference layer undoes these transformations before applying the “right” inference layer $\bm{\mathrm{\beta}}^{*}$ (Eq.˜5, right). In particular, Eq.˜5 (left) guarantees disentanglement of the concepts, ensuring that each learned concept corresponds to a distinct ground-truth concept up to an invertible transformation, matching the notion of [41]. A similar equivalence relation was analyzed for continuous representations in energy-based models, including supervised classifiers, by Khemakhem et al. [35]. CBMs satisfying these conditions are equivalent – specifically by the equivalence relation $\sim$, see Section˜C.1 – to the ground-truth pair $(\mathrm{id},\bm{\mathrm{\beta}}^{*})$; see Fig.˜2 (middle) for an illustration. In Lemma˜C.4, we prove that models with the intended semantics yield the same predictions of the ground-truth pair for all $\bm{\mathrm{g}}\in\mathcal{G}$.
@@ 3.2 Theoretical Analysis
All proofs and the definition of $C[\mathcal{G}]$ can be found in Appendix˜C. Intuitively, the first count includes all deterministic pairs $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$ that achieve maximum likelihood on the training set (i.e., the predicted labels $(\bm{\mathrm{\beta}}\circ\bm{\mathrm{\alpha}})(\bm{\mathrm{g}})$ matches the ground-truth one $\bm{\mathrm{\beta}}^{*}(\bm{\mathrm{g}})$ for all ground-truth concepts $\bm{\mathrm{g}}$ appearing in the data), while the second term subtracts those $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$ that possess the intended semantics as per Definition˜3.3. A positive count implies that a CBM trained via maximum likelihood can learn a deterministic JRS. Notice that the count of JRSs also depends on the choice of the concept space $\mathcal{G}$: a large number of fine-grained ground-truth concept may increase the number of JRSs, while coarser-grained concepts that are lower in number may reduce them.
@@
Corollary˜3.7 implies that when $|\mathsf{Vert}(\mathcal{B})|>1$, the number of deterministic JRSs in Eq.˜6 can be much larger than the number of deterministic RSs (Eq.˜7). For example, in MNIST-SumParity with digits restricted to the range $[0,4]$ there exist $64$ RSs but about $100$ thousand JRSs, and the gap increases as we extend the range $[0,N]$. An in-depth analysis appears in Section˜B.5.
@@
Intuitively, a CBM satisfies this assumption if its inference layer $\bm{\mathrm{\omega}}$ is “peaked”: for any two concept vectors $\bm{\mathrm{c}}$ and $\bm{\mathrm{c}}^{\prime}$ yielding distinct predictions, the label probability output by $\bm{\mathrm{\omega}}$ for any mixture distribution thereof is no larger than the label probability that it associates to $\bm{\mathrm{c}}$ and $\bm{\mathrm{c}}^{\prime}$. That is, distributing probability mass across concepts does not increase label likelihood. While this assumption does not hold for general CBMs, we show in Appendix˜E that it holds for popular architectures, including most of those that we experiment with. Under Assumption˜3.8, we have:
@@ 4 Practical Solutions
By Theorem˜3.9, getting rid of deterministic JRSs from a task prevents CBMs trained via maximum log-likelihood to learn JRS altogether. The key question is how to make the JRSs count be zero. As previously explored in RSs, pairing the maximum likelihood objective with other well-known penalties in the literature can limit optimal $\bm{\mathrm{\alpha}}$’s for the joint training objective. Several mitigation strategies for RSs have shown to decrease the count of the $\bm{\mathrm{\alpha}}$’s, in some cases leading to zeroing deterministic RSs [56, 91]. We report the count reduction of different mitigations in Appendix˜D.
@@ 5 Case Studies
We tackle the following key research questions: Q1. Are CBMs affected by JRSs in practice? Q2. Do JRSs affect interpretability and OOD behavior? Q3. Can existing mitigation strategies prevent JRSs? Appendix˜A reports additional details about the tasks, architectures, and model selection.
Models. We evaluate several (also NeSy) CBMs. DeepProbLog (DPL) [50, 51] is the only method supplied with the ground-truth knowledge $\mathsf{K}$, and uses probabilistic-logic reasoning to ensure predictions are consistent with it. CBNM is a Concept-Bottleneck Model [39] with no supervision on the concepts. Deep Symbolic Learning (DSL) [13] is a SOTA NeSy-CBM that learns concepts and symbolic knowledge jointly; it implements the latter as a truth table and reasoning using fuzzy logic. We also consider two variants: DPL replaces fuzzy with probabilistic logic, while bears wraps DPL within BEARS [53], an ensemble approach for handling regular RSs. In short, bears consists of a learned inference layer and multiple concept extractors. The former is learned only once, the latter are learned sequentially. We also evaluate Self-explainable Neural Networks (SENN) [alvarez2018towards], unsupervised CBMs that include a reconstruction penalty. Since they do not satisfy our assumptions, they are useful to empirically assess the generality of our remarks. In our experiments, we train all CBM variants in a joint manner [39] using only label supervision and do not rely on foundation models for gathering annotations [61, 93, 74]. For additional details, see Appendix˜A.
@@
Metrics. For each CBM, we evaluate predicted labels and concepts with the $F_{1}$ score (resp. $F_{1}(Y)$ and $F_{1}(C)$) on the test split. Computing $F_{1}(C)$ requires to first align the predicted concepts to ground-truth annotations. We do so by using the Hungarian algorithm [40] to find a permutation $\pi$ that maximizes the Pearson correlation across concepts. In this way, we directly evaluate disentanglement of the concepts as per Eq.˜5 (left). Concept collapse $\mathsf{Cls}(C)$ quantifies to what extent the concept extractor blends different ground-truth concepts together: high collapse indicates that it predicts fewer concepts than expected. We measure the quality of the learned inference layer $\bm{\mathrm{\beta}}$ as follows: for each input $\bm{\mathrm{x}}$, (1) we reorder the corresponding concept annotations $\bm{\mathrm{g}}$ using $\pi$, (2) feed the result to $\bm{\mathrm{\beta}}$, and (3) compute the $F_{1}$-score ($F_{1}(\bm{\mathrm{\beta}})$) of its predictions. Step (1) permutes and applies element-wise invertible transformations on the ground-truth concepts using the inverse of the map returned by Hungarian matching. This evaluates whether the learned inference layer can correctly handle (an invertible transformation of) the ground-truth concept annotations, measuring the generalization as per Eq.˜5 (right).This procedure is precisely detailed in Section˜A.6.
@@
Q3: Popular CBM and RS mitigations fall short. Finally, we test mitigation strategies on CBNM trained on MNIST-SumParity and Clevr, ablating the amount of supervision. The results for MNIST-SumParity in Fig.˜3 show while concept supervision ($x$-axis) helps all metrics, adding reconstruction, entropy maximization, and contrastive learning to the mix brings no benefits for concept and knowledge ($F_{1}(C)$, $F_{1}(\bm{\mathrm{\beta}})$) quality, not even when combined (orange curve). Knowledge supervision is also ineffective (blue lines), likely because MNIST-SumParity and Clevr suffer from RSs even when the model is supplied prior knowledge. Contrastive learning improves concept collapse ($\mathsf{Cls}(C)$): the orange and purple curves in Fig.˜3 (right) show it can prevent CBMs from mixing concepts together. The results for Clevr in Appendix˜B show similar trends.

### v4

@@ 1 Introduction
This raises the question of when CBMs can acquire “high-quality” concepts and inference layers. While existing studies focus on concept quality [49, 99, 55], they neglect the role of the inference layer altogether. In contrast, we cast the question in terms of whether it is possible to identify from data concepts and inference layers with the intended semantics, defined formally in Section˜3.1. We proceed to answer this question by building a novel connection with reasoning shortcuts (RSs), a well-known issue in Neuro-Symbolic (NeSy) AI whereby models achieve high accuracy by learning low-quality concepts even if the inference layer is fixed [9, 56, 84, 91]. For instance, a NeSy model for autonomous driving whose inference layer encodes the traffic laws might confuse pedestrians with red lights, as both entail the same prediction (the car has to stop) [4]. We generalize RSs to CBMs in which the inference layer is learned, and concept supervision may be absent. Our analysis shows that shortcuts can be exponentially many, even more than RSs, (we count them explicitly in Section˜9.5) and that maximum likelihood training is insufficient for attaining intended semantics. This hinders both interpretability and OOD behavior. On the positive side, we also specify conditions under which (under suitable assumptions) CBMs cannot be fooled by RSs, proving that the ground-truth concepts and inference layer can be identified (see Theorem˜3.9).
@@ 3 Reasoning Shortcuts in CBMs
Here, $\bm{\mathrm{\omega}}_{\bm{\mathrm{y}}}$ is the conditional probability of the ground-truth labels $\bm{\mathrm{y}}$. Notice that under the above assumptions CBMs attaining maximum likelihood perfectly model the ground-truth data distribution $p^{*}(\bm{\mathrm{X}},\bm{\mathrm{Y}})$, see Lemma˜10.1.
@@ 3.1 Intended Semantics and Joint Reasoning Shortcuts
Intended semantics holds if the learned concepts $\bm{\mathrm{C}}$ match the ground-truth concepts $\bm{\mathrm{G}}$ modulo simple invertible transformations – like reordering and negation (Eq.˜5, left) – and the inference layer undoes these transformations before applying the “right” inference layer $\bm{\mathrm{\beta}}^{*}$ (Eq.˜5, right). In particular, Eq.˜5 (left) guarantees disentanglement of the concepts, ensuring that each learned concept corresponds to a distinct ground-truth concept up to an invertible transformation, matching the notion of [41]. A similar equivalence relation was analyzed for continuous representations in energy-based models, including supervised classifiers, by Khemakhem et al. [35]. CBMs satisfying these conditions are equivalent – specifically by the equivalence relation $\sim$, see Section˜10.1 – to the ground-truth pair $(\mathrm{id},\bm{\mathrm{\beta}}^{*})$; see Fig.˜2 (middle) for an illustration. In Lemma˜10.4, we prove that models with the intended semantics yield the same predictions of the ground-truth pair for all $\bm{\mathrm{g}}\in\mathcal{G}$.
@@ 3.2 Theoretical Analysis
All proofs and the definition of $C[\mathcal{G}]$ can be found in Section˜10. Intuitively, the first count includes all deterministic pairs $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$ that achieve maximum likelihood on the training set (i.e., the predicted labels $(\bm{\mathrm{\beta}}\circ\bm{\mathrm{\alpha}})(\bm{\mathrm{g}})$ matches the ground-truth one $\bm{\mathrm{\beta}}^{*}(\bm{\mathrm{g}})$ for all ground-truth concepts $\bm{\mathrm{g}}$ appearing in the data), while the second term subtracts those $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$ that possess the intended semantics as per Definition˜3.3. A positive count implies that a CBM trained via maximum likelihood can learn a deterministic JRS. Notice that the count of JRSs also depends on the choice of the concept space $\mathcal{G}$: a large number of fine-grained ground-truth concept may increase the number of JRSs, while coarser-grained concepts that are lower in number may reduce them.
@@
Corollary˜3.7 implies that when $|\mathsf{Vert}(\mathcal{B})|>1$, the number of deterministic JRSs in Eq.˜6 can be much larger than the number of deterministic RSs (Eq.˜7). For example, in MNIST-SumParity with digits restricted to the range $[0,4]$ there exist $64$ RSs but about $100$ thousand JRSs, and the gap increases as we extend the range $[0,N]$. An in-depth analysis appears in Section˜9.5.
@@
Intuitively, a CBM satisfies this assumption if its inference layer $\bm{\mathrm{\omega}}$ is “peaked”: for any two concept vectors $\bm{\mathrm{c}}$ and $\bm{\mathrm{c}}^{\prime}$ yielding distinct predictions, the label probability output by $\bm{\mathrm{\omega}}$ for any mixture distribution thereof is no larger than the label probability that it associates to $\bm{\mathrm{c}}$ and $\bm{\mathrm{c}}^{\prime}$. That is, distributing probability mass across concepts does not increase label likelihood. While this assumption does not hold for general CBMs, we show in Section˜12 that it holds for popular architectures, including most of those that we experiment with. Under Assumption˜3.8, we have:
@@ 4 Practical Solutions
By Theorem˜3.9, getting rid of deterministic JRSs from a task prevents CBMs trained via maximum log-likelihood to learn JRS altogether. The key question is how to make the JRSs count be zero. As previously explored in RSs, pairing the maximum likelihood objective with other well-known penalties in the literature can limit optimal $\bm{\mathrm{\alpha}}$’s for the joint training objective. Several mitigation strategies for RSs have shown to decrease the count of the $\bm{\mathrm{\alpha}}$’s, in some cases leading to zeroing deterministic RSs [56, 91]. We report the count reduction of different mitigations in Section˜11.
@@ 5 Case Studies
We tackle the following key research questions: Q1. Are CBMs affected by JRSs in practice? Q2. Do JRSs affect interpretability and OOD behavior? Q3. Can existing mitigation strategies prevent JRSs? Section˜8 reports additional details about the tasks, architectures, and model selection.
Models. We evaluate several (also NeSy) CBMs. DeepProbLog (DPL) [50, 51] is the only method supplied with the ground-truth knowledge $\mathsf{K}$, and uses probabilistic-logic reasoning to ensure predictions are consistent with it. CBNM is a Concept-Bottleneck Model [39] with no supervision on the concepts. Deep Symbolic Learning (DSL) [13] is a SOTA NeSy-CBM that learns concepts and symbolic knowledge jointly; it implements the latter as a truth table and reasoning using fuzzy logic. We also consider two variants: DPL replaces fuzzy with probabilistic logic, while bears wraps DPL within BEARS [53], an ensemble approach for handling regular RSs. In short, bears consists of a learned inference layer and multiple concept extractors. The former is learned only once, the latter are learned sequentially. We also evaluate Self-explainable Neural Networks (SENN) [alvarez2018towards], unsupervised CBMs that include a reconstruction penalty. Since they do not satisfy our assumptions, they are useful to empirically assess the generality of our remarks. In our experiments, we train all CBM variants in a joint manner [39] using only label supervision and do not rely on foundation models for gathering annotations [61, 93, 74]. For additional details, see Section˜8.
@@
Metrics. For each CBM, we evaluate predicted labels and concepts with the $F_{1}$ score (resp. $F_{1}(Y)$ and $F_{1}(C)$) on the test split. Computing $F_{1}(C)$ requires to first align the predicted concepts to ground-truth annotations. We do so by using the Hungarian algorithm [40] to find a permutation $\pi$ that maximizes the Pearson correlation across concepts. In this way, we directly evaluate disentanglement of the concepts as per Eq.˜5 (left). Concept collapse $\mathsf{Cls}(C)$ quantifies to what extent the concept extractor blends different ground-truth concepts together: high collapse indicates that it predicts fewer concepts than expected. We measure the quality of the learned inference layer $\bm{\mathrm{\beta}}$ as follows: for each input $\bm{\mathrm{x}}$, (1) we reorder the corresponding concept annotations $\bm{\mathrm{g}}$ using $\pi$, (2) feed the result to $\bm{\mathrm{\beta}}$, and (3) compute the $F_{1}$-score ($F_{1}(\bm{\mathrm{\beta}})$) of its predictions. Step (1) permutes and applies element-wise invertible transformations on the ground-truth concepts using the inverse of the map returned by Hungarian matching. This evaluates whether the learned inference layer can correctly handle (an invertible transformation of) the ground-truth concept annotations, measuring the generalization as per Eq.˜5 (right).This procedure is precisely detailed in Section˜8.6.
@@
Q3: Popular CBM and RS mitigations fall short. Finally, we test mitigation strategies on CBNM trained on MNIST-SumParity and Clevr, ablating the amount of supervision. The results for MNIST-SumParity in Fig.˜3 show while concept supervision ($x$-axis) helps all metrics, adding reconstruction, entropy maximization, and contrastive learning to the mix brings no benefits for concept and knowledge ($F_{1}(C)$, $F_{1}(\bm{\mathrm{\beta}})$) quality, not even when combined (orange curve). Knowledge supervision is also ineffective (blue lines), likely because MNIST-SumParity and Clevr suffer from RSs even when the model is supplied prior knowledge. Contrastive learning improves concept collapse ($\mathsf{Cls}(C)$): the orange and purple curves in Fig.˜3 (right) show it can prevent CBMs from mixing concepts together. The results for Clevr in Section˜9 show similar trends.